OpenStudy (anonymous):

is this TRUE or FALSE (x-9)(x+4)=xsquared-5x-36

OpenStudy (anonymous):

True statement

OpenStudy (anonymous):

how ? to solve for it??

OpenStudy (anonymous):

Multiply the terms inside the parentheses with each other \[(x-9)(x+4)\] \[=x^2+4x-9x-36\] \[=x^2-5x-36\]

OpenStudy (anonymous):

oh ayh thnx iforgot the 9 was a negative i added it thnx!

OpenStudy (anonymous):

thanks to you i know whats my mistake